[ [Genny] & Jill left home at the same time. They traveled in opposite directions. Jill traveled 15 mph faster than [Genny]. After 6 hrs. they were 630 km apart. Find [Genny's] speed. ]
Answer by TimothyLamb(4379)

g = genny's speed
j = jill's speed
---
j = g + 15
---
since they travel in exactly opposite directions, the vector math for their relative speed is trivial, and we can simplify by assuming a single traveler moving at the relative speed of the two travelers.
---
relative speed:
s = g + (g + 15)
s = 2g + 15
---
s = d / t
s = 630/6
2g + 15 = 630/6
2g + 15 = 105
2g = 105 - 15
2g = 90
---
g = 45 mph
j = 60 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php